write an inequality on the variable x for the degree measure of the smallest ngle of the triangle , given tha the gegree measure ofthe smallest angle is at most 30 degrees.
Answer by stanbon(75887)

write an inequality on the variable x for the degree measure of the smallest ngle of the triangle , given tha the gegree measure ofthe smallest angle is at most 30 degrees.
---
x + y = 180
y = 180-x
---
Since x <= 30, y >= 150
-------
Therefore:
180-x >= 150
=====================
Cheers,
Stan H.